A ladder is leaning against a vertical wall, and both ends of the ladder are at the point of slipping. The coefficient of friction between the ladder and the horizontal surface is μ1 = 0.275 and the coefficient of friction between the ladder and the wall is μ2 = 0.153. Determine the maximum angle with the vertical the ladder can make without falling on the ground.
Answer #1

W=ladder weight
Rx=floor friction force
Ry=floor normal reaction
H=wall normal reaction
V=wall friction force [V=μ2H]
sum of forces must be null
then
H=Rx=μ1W [horizontal forces]
V=W-Ry or μ2H=W-Ry or μ1μ2W=W-Ry [vertical forces] (*)
sum of torques must be null
then [assuming the wall contact point as pole]
RyLcosθ=W(L/2)sinθ
Ry=W/(2tanθ)
returning to eq.(*) on get
μ1μ2W=W-W/(2tanθ)
from which
tanθ=1/[2(1-μ1μ2)]
tanθ=1/[2(1-0.275*0.153)] = 0.52196
θ=27.56 degrees